OpenStudy (anonymous):

When a mortar shell is fired with an initial velocity of v0 ft/sec at an angle α above the horizontal, then its position after t seconds is given by the parametric equations x = (v0 cos α)t , y = (v0 sin α)t − 16t^2. If the mortar shell hits the ground 2500 feet from the mortar when α = 75degress, determine v0 *answer choices* 1. v0 = 360 ft/sec 2. v0 = 370 ft/sec 3. v0 = 380 ft/sec 4. v0 = 390 ft/sec 5. v0 = 400 ft/sec

OpenStudy (experimentx):

There must be some error.

OpenStudy (anonymous):

exactly 400m/s

OpenStudy (anonymous):

\[u ^{2}=Rg/\sin2\alpha\] solving for u gives u =\[\sqrt{160000}=400ft/s\] remember g=32ft/s

OpenStudy (anonymous):

\[Range(R)=u ^{2}\sin2\alpha/g\] here range=2500ft \[\alpha\]=75

OpenStudy (experimentx):

must be right.